OpenStudy (anonymous):

Two balls of mass = 1.00 kg are floating 10 feet apart. What is the nature of the gravitational forces these balls exert on each other?

OpenStudy (abhisar):

Hey there! Gravitational forces are always attractive forces and hence will tend to bring the two balls closer to each other. You can also calculate the magnitude of the gravitational force that the two balls will exert on each other by using the following equation. \(\huge \boxed{\tt F_g=\frac{G\times m_1\times m_2}{r^2}}\) Here, \(\bullet ~\tt F_g\) is the gravitational force. \(\bullet~\) G is the universal gravitational constant = \(\sf 6.67384 × 10^{-11} m^3 kg^{-1} s^{-2}\) \(\bullet~\)\( \tt m_1~and~m_2\) are the masses of two balls respectively. \(\bullet~\)r is the distance between the two balls

OpenStudy (anonymous):

a. Fball 1 on ball 2 = Fball 2 on ball 1 b. Fball 1 on ball 2 = -Fball 2 on ball 1 c. Fball 1 on ball 2 < Fball 2 on ball 1 d. Fball 1 on ball 2 > Fball 2 on ball 1 ok, but how does it relate to the answers?